An endothermic chemical reaction is conducted in a batch reactor with a heating jacket. A feedstock of 150 Kg is loaded in the reactor at 20o C and the reaction proceeds for 1 hour. At the end of the process, products are at 100o C.. The reaction absorbs 2300 kJ/kg of material in the reactor. Heat is provided by saturated steam. The steam enters the heating jacket at 250oC and fully condenses. The systems has heat losses equal to 1.5 kJ/h.

Reagents and products have an average heat capacity of Cp=3.26J/kg.

a) Draw a flowsheet to represent this process

b) Calculate the amount of steam needed per kg of reagents
Answer #1

a)

Sits

Q is the heat lost to the surrounding

b) saturated steam at 250 c is used then

from steam stable we calculate enthalpy of steam at 250c

= 1716.2 KJ/kg

So S is the kg of steam used then (S*1716.2 )KJ of heat provided by the steam

And this is equal to heat absorb by the reaction + heat lost to surrounding

S*1716.2 = (2300*150)+(1.5*1)

S = 201 Kg

This is the amount of steam require for 150 kg of reactant

Then steam required per kg of reactant = 201/150 = 1.34 kg